Herald, a barista, earned 260,000 in 2019. But in 2021, he was retrenched and decided to open his own coffee shop. After a year, he submitted the following data :

  •   He spent 45,000 on equipment.
  •   He leased a small space for P70,000 a year.
  •   He paid P 100,000 in wages.
  •   He sold worth P950,000 of coffee a year.
  •   He used P150,000 from his saving account, which pays 3% interest a year. 
  •   He borrowed P250,000 at 10% a year from the bank.
  •   Normal profit is P350,000 a year.

 

Compute for the following :

  •    Herald's explicit costs (10 points)
  •    Herald's implicit cost (10 points)
  •    Herald's economic profit ( 10 points)
1
Expert's answer
2021-11-22T09:56:46-0500

Solution:

1.). Explicit costs:

Purchase of equipment = 45,000

Rent = 70,000

Wages = 100,000

Loan bank interest = 10"\% \times" 250,000 = 25,000

Explicit costs = 45,000 + 70,000 + 100,000 + 25,000 = 240,000

 

2.). Implicit costs:

Heralds previous earning = 260,000

Savings account interest = 150,000 "\times3\%" = 4,500

Total implicit costs = 260,000 + 4,500 = 264,500

 

3.). Economic profit:

Economic profit = Total revenue – (Explicit costs + Implicit costs)

Economic profit = 950,000 – (240,000 + 264,500) = 950,000 – 504,500 = 445,500

Economic profit = 445,500